OpenStudy (anonymous):

Not quite sure if I understand this. Use the Pythagorean Theorem to find the length of the missing side of the right triangle with both legs equal to √2.

OpenStudy (mathstudent55):

Do you know the Pythagorean Theorem?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I understand how it works when the given sides of the triangle are not square roots

OpenStudy (anonymous):

The radicals are throwing me off

OpenStudy (mathstudent55):

It works the same even if the sides are square roots. \(a^2 + b^2 = c^2\) a and b are the lengths of the legs. c is the length of the hypotenuse. You are told that both a and b are \(\sqrt 2\). Just replace a and b in the equation with \(\sqrt 2\). What do you get?